With the wind, a jet takes 3 hours to fly 1,890 miles. Against the wind, it takes 3 3/8 hours for the same trip. Find the wind speed and the plane's air speed.
If x is the plane's speed and y is the wind's speed, then which two expressions represent the rate of speed in each direction?

Answer by scott8148(6628)

x + y = 1890 / 3

x - y = 1890 / (27/8)

add the equations to find x (y cancels), then substitute to find y